# Interplay of correlations and Majorana mode from local solution perspective

### v1

@@ I Introduction
It has been established [33] that charge tunneling could probe the lifetime of the Majorana states in heterostructures consisting of a metal-quantum dot-topological superconductor. Charge transfer varies the electron number on the quantum dot by ±$\pm 1$, thus connecting the even and odd parity sections. In what follows, we determine the probability of such parity changes in the strongly correlated quantum dot. This brings information concerning optimal conditions for the leakage of the Majorana modes.
@@ II Eigenstates and eigenenergies
It is convenient to express the Majorana operators in terms of the conventional fermion operators $\hat{f},\hat{f^{\dagger}}$ defined through $\hat{\eta}_{1}=\frac{1}{\sqrt{2}}(\hat{f}^{\dagger}+\hat{f})$ and $\hat{\eta}_{2}=\frac{i}{\sqrt{2}}(\hat{f}^{\dagger}-\hat{f})$. Hamiltonian (1) then acquires the following structure:
$\hat{H}$ $=$ $\hat{H}_{QD}+t_{m}(\hat{d}_{\downarrow}^{\dagger}\hat{f}+\hat{f}^{\dagger}\hat{d}_{\downarrow})$ (3)
@@
[…]
@@ III.1 Spectrum of spin-$\downarrow$ electrons
On the other hand, transitions between the states $\Psi^{s}_{1}\leftrightarrow\Psi^{\bar{s}}_{2}$ and $\Psi^{s}_{3}\leftrightarrow\Psi^{\bar{s}}_{4}$ with different sign index $s\neq\bar{s}$ contribute to the finite-energy poles at ±$\pm E_{p}$ and ±$\pm E_{q}$, respectively. For $\Psi^{s}_{1}\leftrightarrow\Psi^{\bar{s}}_{2}$ matrix elements are given by $|\langle\Psi_{1}^{+}|d_{\downarrow}|\Psi_{2}^{-}\rangle|^{2}=|\langle\Psi_{2}^{+}|d_{\downarrow}|\Psi_{1}^{-}\rangle|^{2}=v_{p}^{4}$ and $|\langle\Psi_{1}^{-}|d_{\downarrow}|\Psi_{2}^{+}\rangle|^{2}=|\langle\Psi_{2}^{-}|d_{\downarrow}|\Psi_{1}^{+}\rangle|^{2}=u_{p}^{4}$. The contribution to the Green’s function from the first two transitions takes the form:
@@
Figure 3 shows the typical spectrum of $\downarrow$-spin electrons. The black dashed line indicates the zero-energy quasiparticle, originating from the Majorana mode leakage. Red/green dashed lines correspond to the quasiparticle energies ±$\pm E_{p}$ and blue/magenta indicate the quasiparticle energies ±$\pm E_{q}$, respectively. To understand their physical meaning, let us recall that an isolated QD ($t_{m}=0$) has two quasiparticle energies: at $\omega=\varepsilon_{d}$ (i.e. $\xi_{d}=-\frac{U_{d}}{2}$) with spectral weight $1-n_{d\sigma}$ and another Coulomb satellite at $\omega=\varepsilon_{d}+U_{d}$ (i.e. $\xi_{d}=\frac{U_{d}}{2}$) with spectral weight $n_{d\sigma}$. For $t_{m}\neq 0$, these quasiparticle branches evolve into the trivial modes ±$\pm E_{p}$ and ±$\pm E_{q}$ of our setup, which are gapped due to the intersite pairing (for details see Sec. III.4). In Figure 3, we clearly notice avoided-crossing behavior of the trivial (finite-energy) quasiparticle branches, repelled at some distance from the topological (zero-energy) mode, which is due to the protection of the Majorana state. At the critical points, $\xi_{d}=\pm\frac{U_{d}}{2}$, the trivial states are separated from the zero-energy mode by a gap of $2t_{m}$. Furthermore, the spectral weight $A_{1}$ of the zero-energy mode, $\omega=0$, is enhanced around $\omega=\xi_{d}-\frac{U_{d}}{2}$ and $\omega=\xi_{d}+\frac{U_{d}}{2}$.
@@ III.3 Magnetization
[…]
@@ III.4 Signatures of intersite pairing
Figure 9: Transmittance of the selective equal spin Andreev reflection (SESAR) obtained for $t_{m}/U_{d}=0.25$ (top panel), $0.5$ (middle panel), and $1$ (bottom panel), assuming temperature $k_{b}T=0.01U_{d}$ and coupling $\Gamma_{N}=0.01U_{d}$.
@@
±$\pm\frac{1}{Z}\frac{2t_{m}}{E_{p}}(e^{-\beta E_{1}^{-}}+e^{-\beta E_{1}^{+}})\frac{1}{\omega+i\Gamma_{N}\pm E_{p}},$ (53)
±$\pm\frac{1}{Z}\frac{2t_{m}}{E_{q}}(e^{-\beta E_{3}^{-}}+e^{-\beta E_{3}^{+}})\frac{1}{\omega+i\Gamma_{N}\pm E_{q}}.$ (54)
@@ IV Results for $\epsilon_{m}\neq 0$
$\rho_{\downarrow}(\omega)$ $=$ $A_{12}^{-/-}\delta\left[\omega+\frac{1}{2}(E^{-}_{p}-E^{+}_{p})\right]$ (61)
$+$ $A_{12}^{-/+}\delta\left[\omega-\frac{1}{2}(E^{-}_{p}+E^{+}_{p})\right]$
$+$ $A_{12}^{+/-}\delta\left[\omega+\frac{1}{2}(E^{-}_{p}+E^{+}_{p})\right]$
$+$ $A_{12}^{+/+}\delta\left[\omega-\frac{1}{2}(E^{-}_{p}-E^{+}_{p})\right]$
$+$ $A_{34}^{-/-}\delta\left[\omega-\frac{1}{2}(E^{-}_{q}-E^{+}_{q})\right]$
$+$ $A_{34}^{-/+}\delta\left[\omega-\frac{1}{2}(E^{-}_{q}+E^{+}_{q})\right]$
$+$ $A_{34}^{+/-}\delta\left[\omega+\frac{1}{2}(E^{-}_{q}+E^{+}_{q})\right]$
$+$ $A_{34}^{+/+}\delta\left[\omega+\frac{1}{2}(E^{-}_{q}-E^{+}_{q})\right]$
@@
$+$ $\frac{1}{Z}(v_{1}^{+}v_{2}^{+})^{2}(e^{-\beta E^{+}_{1}}+e^{-\beta E^{+}_{2}}),$
$A^{-/+}_{12}$ $=$ $\frac{1}{Z}(u_{1}^{-}u_{2}^{+})^{2}(e^{-\beta E^{-}_{1}}+e^{-\beta E^{+}_{2}})$
$+$ $\frac{1}{Z}(v_{1}^{-}v_{2}^{+})^{2}(e^{-\beta E^{-}_{1}}+e^{-\beta E^{+}_{2}}),$
@@
$+$ $\frac{1}{Z}(v_{3}^{+}v_{4}^{+})^{2}(e^{-\beta E^{+}_{3}}+e^{-\beta E^{+}_{4}}),$
$A^{-/+}_{34}$ $=$ $\frac{1}{Z}(u_{3}^{-}u_{4}^{+})^{2}(e^{-\beta E^{-}_{3}}+e^{-\beta E^{+}_{4}})$
$+$ $\frac{1}{Z}(v_{3}^{-}v_{4}^{+})^{2}(e^{-\beta E^{-}_{3}}+e^{-\beta E^{+}_{4}}),$
@@ V Summary and outlook
For the non-correlated case, the trivial quasiparticles exist at energies ±$\pm\sqrt{\varepsilon_{d}^{2}+4t_{m}^{2}}$. In this scenario, the optimal amplitude of the zero-energy mode occurs for $\varepsilon_{d}=0$. Under such circumstances, the Majorana mode acquires half of the total spectral weight, and the trivial quasiparticles equally share the remaining amount. The spectral function of QD in this case is represented by a three-peak structure $\rho_{\downarrow}(\omega)=0.5\delta(\omega)+0.25\delta(\omega-2t_{m})+0.25\delta(\omega+2t_{m})$. Away from half-filling, one of the trivial quasiparticles gradually absorbs more and more spectral weight, at the expense of both the other conventional quasiparticle and the zero-energy mode.
In the presence of the Coulomb repulsion, a leakage of the zero-energy mode is most efficient when the zero mode coincides either with the energy level $\varepsilon_{d}=0$ or with the Coulomb satellite $\varepsilon_{d}+U_{d}=0$ (i.e. $\xi_{d}=\pm\frac{U_{d}}{2}$). One should note that, at such points, the trivial quasiparticles are formed at ±$\pm 2t_{m}$, provided that the Majorana modes do not overlap with one another. Away from these points, the spectrum of $\downarrow$-spin electrons consists of four trivial quasiparticles coexisting with the zero-energy mode (Fig.3). We have demonstrated that they could be experimentally detected by spin-polarized Andreev spectroscopy, Fig. 9. The spin-$\uparrow$ sector also consists of four quasiparticle branches, but all of them refer to the nontopological states.
@@ Appendix A Influence of Zeeman field
Figure 13: Dependence of the eigenenergies $E_{i}^{-}$ on the energy level $\varepsilon_{d}$ of QD in Zeeman field $h=0.5U_{d}$ obtained for weak overlap $\epsilon_{m}=0.1U_{d}$

### v2

@@ I Introduction
It has been established [33] that charge tunneling could probe the lifetime of the Majorana states in heterostructures consisting of a metal-quantum dot-topological superconductor. Charge transfer varies the electron number on the quantum dot by $\pm 1$, thus connecting the even and odd parity sections. In what follows, we determine the probability of such parity changes in the strongly correlated quantum dot. This brings information concerning optimal conditions for the leakage of the Majorana modes.
@@ II Eigenstates and eigenenergies
[…]
@@
Let us remark that the correlated quantum dot coupled to a conventional superconductor would be characterized by a different set of eigenvectors, represented either by the singly occupied configurations $|\uparrow\rangle$ and $|\downarrow\rangle$ or the BCS-type coherent superpositions $u^{\pm}|0\rangle+v^{\pm}|\uparrow\downarrow\rangle$ [34, 35]. Here, in contrast, we obtain eigenstates that are superpositions of either the empty and singly occupied dot, $|\Psi_{1,2}\rangle$, or the doubly and singly occupied dot, $|\Psi_{3,4}\rangle$, combined with the edge mode. Unlike the mentioned BCS-type superpositions, the eigenstates of the considered system are superpositions of states with different dot electron parity. One can also note that the states $|\Psi_{1,2}\rangle$ are characterized by opposite dot magnetization compared to states $|\Psi_{3,4}\rangle$. Consequently, a ground state transition from $|\Psi_{1,2}\rangle$ to $|\Psi_{3,4}\rangle$ (or vice versa) is accompanied by a conversion of the dot’s magnetic properties (c.f. Fig. 2). Such a set of eigenfunctions originates from the intersite pairing. For each of these configurations, we obtained two possible solutions, $\hat{H}|\Psi_{i}^{\pm}\rangle=E_{i}^{\pm}|\Psi_{i}^{\pm}\rangle$, with eigenvalues
@@ III.1 Spectrum of spin-$\downarrow$ electrons
On the other hand, transitions between the states $\Psi^{s}_{1}\leftrightarrow\Psi^{\bar{s}}_{2}$ and $\Psi^{s}_{3}\leftrightarrow\Psi^{\bar{s}}_{4}$ with different sign index $s\neq\bar{s}$ contribute to the finite-energy poles at $\pm E_{p}$ and $\pm E_{q}$, respectively. For $\Psi^{s}_{1}\leftrightarrow\Psi^{\bar{s}}_{2}$ matrix elements are given by $|\langle\Psi_{1}^{+}|d_{\downarrow}|\Psi_{2}^{-}\rangle|^{2}=|\langle\Psi_{2}^{+}|d_{\downarrow}|\Psi_{1}^{-}\rangle|^{2}=v_{p}^{4}$ and $|\langle\Psi_{1}^{-}|d_{\downarrow}|\Psi_{2}^{+}\rangle|^{2}=|\langle\Psi_{2}^{-}|d_{\downarrow}|\Psi_{1}^{+}\rangle|^{2}=u_{p}^{4}$. The contribution to the Green’s function from the first two transitions takes the form:
@@
Figure 3 shows the typical spectrum of $\downarrow$-spin electrons. The black dashed line indicates the zero-energy quasiparticle, originating from the Majorana mode leakage. Red/green dashed lines correspond to the quasiparticle energies $\pm E_{p}$ and blue/magenta indicate the quasiparticle energies $\pm E_{q}$, respectively. To understand their physical meaning, let us recall that an isolated QD ($t_{m}=0$) has two quasiparticle energies: at $\omega=\varepsilon_{d}$ (i.e. $\xi_{d}=-\frac{U_{d}}{2}$) with spectral weight $1-n_{d\sigma}$ and another Coulomb satellite at $\omega=\varepsilon_{d}+U_{d}$ (i.e. $\xi_{d}=\frac{U_{d}}{2}$) with spectral weight $n_{d\sigma}$. For $t_{m}\neq 0$, these quasiparticle branches evolve into the trivial modes $\pm E_{p}$ and $\pm E_{q}$ of our setup, which are gapped due to the intersite pairing (for details see Sec. III.4). In Figure 3, we clearly notice avoided-crossing behavior of the trivial (finite-energy) quasiparticle branches, repelled at some distance from the topological (zero-energy) mode, which is due to the protection of the Majorana state. At the critical points, $\xi_{d}=\pm\frac{U_{d}}{2}$, the trivial states are separated from the zero-energy mode by a gap of $2t_{m}$. Furthermore, the spectral weight $A_{1}$ of the zero-energy mode, $\omega=0$, is enhanced around $\omega=\xi_{d}-\frac{U_{d}}{2}$ and $\omega=\xi_{d}+\frac{U_{d}}{2}$.
@@ III.3 Magnetization
emerging outside the half-filling (see Fig. 7). To explain the sign change of QD magnetization, let us inspect Eqs. (4-7), noting that $\Psi_{1}$ and $\Psi_{2}$ represent superpositions of the empty and singly occupied spin-$\downarrow$ configurations. Therefore, if for particular model parameters state $\Psi_{1}$ or $\Psi_{2}$ is the ground state, the dot magnetization would be oriented along the $\downarrow$-direction. In contrast, the states $\Psi_{3}$ and $\Psi_{4}$ are superpositions of the nonmagnetic $|\uparrow\downarrow\rangle$ state and the singly occupied spin-$\uparrow$ configuration. The ground state represented by $\Psi_{3}$ and $\Psi_{4}$ would then have magnetization along the $\uparrow$-direction. Fig. 2 shows that for $\xi_{d}<0$, the ground state of QD is represented by $\Psi_{3}$ or $\Psi_{4}$. This fact explains the sign change of the magnetization at half-filling, $\xi_{d}=0$, in agreement with previous studies reported in Refs.[10, 11, 36].
@@ III.4 Signatures of intersite pairing
Figure 9: Transmittance of the selective equal spin Andreev reflection (SESAR) obtained for $t_{m}/U_{d}=0.25$ (top panel), $0.5$ (middle panel), and 11 (bottom panel), assuming temperature $k_{b}T=0.01U_{d}$ and coupling $\Gamma_{N}=0.01U_{d}$.
@@
$\pm\frac{1}{Z}\frac{2t_{m}}{E_{p}}(e^{-\beta E_{1}^{-}}+e^{-\beta E_{1}^{+}})\frac{1}{\omega+i\Gamma_{N}\pm E_{p}},$ (53)
$\pm\frac{1}{Z}\frac{2t_{m}}{E_{q}}(e^{-\beta E_{3}^{-}}+e^{-\beta E_{3}^{+}})\frac{1}{\omega+i\Gamma_{N}\pm E_{q}}.$ (54)
@@ IV Results for $\epsilon_{m}\neq 0$
[…]
@@
[…]
@@
[…]
@@ V Summary and outlook
For the non-correlated case, the trivial quasiparticles exist at energies $\pm\sqrt{\varepsilon_{d}^{2}+4t_{m}^{2}}$. In this scenario, the optimal amplitude of the zero-energy mode occurs for $\varepsilon_{d}=0$. Under such circumstances, the Majorana mode acquires half of the total spectral weight, and the trivial quasiparticles equally share the remaining amount. The spectral function of QD in this case is represented by a three-peak structure $\rho_{\downarrow}(\omega)=0.5\delta(\omega)+0.25\delta(\omega-2t_{m})+0.25\delta(\omega+2t_{m})$. Away from half-filling, one of the trivial quasiparticles gradually absorbs more and more spectral weight, at the expense of both the other conventional quasiparticle and the zero-energy mode.
In the presence of the Coulomb repulsion, a leakage of the zero-energy mode is most efficient when the zero mode coincides either with the energy level $\varepsilon_{d}=0$ or with the Coulomb satellite $\varepsilon_{d}+U_{d}=0$ (i.e. $\xi_{d}=\pm\frac{U_{d}}{2}$). One should note that, at such points, the trivial quasiparticles are formed at $\pm 2t_{m}$, provided that the Majorana modes do not overlap with one another. Away from these points, the spectrum of $\downarrow$-spin electrons consists of four trivial quasiparticles coexisting with the zero-energy mode (Fig.3). We have demonstrated that they could be experimentally detected by spin-polarized Andreev spectroscopy, Fig. 9. The spin-$\uparrow$ sector also consists of four quasiparticle branches, but all of them refer to the nontopological states.
@@ Appendix A Influence of Zeeman field
Figure 13: Dependence of the eigenenergies $E_{i}^{-}$ on the energy level $\varepsilon_{d}$ of QD in Zeeman field $h=0.5U_{d}$ obtained for weak overlap $\epsilon_{m}=0.1U_{d}$ .